A solution is made by mixing 35.0 mL of ethanol, C2H6O, and 65.0 mL of water.
Assuming ideal behavior, what is the vapor pressure of the solution at 20
Answer #1
moles of C2H6O = (35x.789)/46 = 0.6
moles of H2O = (65x.998)/18 = 3.6

mole fraction of H2O = 3.6/(.6+3.6) = 0.857
mole fraction of C2H6O = 1-.857 = .143

pC2H6O = .143 x 43.9 = 6.27 torr
pH2O = .857 x 17.5 = 14.99 torr

Vapour pressure of soln = 6.27+14.99= 21.26 torr